OpenStudy (anonymous):

If f(x) varies directly with x, and f(x) = 45 when x = 9, find the value of f(x) when x = 3. A. 5 B. 9 C. 15 D. 27

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

f(x) = 45 x = 9 Plug them in: 45 = k(9) 45 = 9k Divide 9 to both sides, what's 45 / 9?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

And your using "K". Does it matter which one you use? Because there is no "K" in the equation

OpenStudy (adi3):

5 is right

OpenStudy (igreen):

Yes, so the constant is '5'. No, it doesn't..it just represents the constant..that's how they use it.

OpenStudy (igreen):

Now let's plug in k = 5 and x = 3 and solve for f(x). f(x) = kx f(x) = (5)(3) Multiply 5 * 3

OpenStudy (anonymous):

Ok. I have a few more. Help me with those also?

OpenStudy (adi3):

no wait the method which teacher was t o45/9 = 5 5 * 3 = 15 answer is 15

OpenStudy (adi3):

no no wait I messed up it was for some other subject not fo this one 5 is right

OpenStudy (anonymous):

:?

OpenStudy (anonymous):

So the answer is 15 @iGreen

OpenStudy (igreen):

Yes, it is.

OpenStudy (adi3):

wht is am right

OpenStudy (adi3):

@iGreen

OpenStudy (anonymous):

Ok. I have a cpiuple more. Help me with those?

OpenStudy (anonymous):

couple*

OpenStudy (adi3):

guys some one give me medal

OpenStudy (adi3):

pls

OpenStudy (anonymous):

I already gave one to @iGreen Sorry

OpenStudy (adi3):

redo it

OpenStudy (adi3):

@DelTa-_-LeGEnD